OpenStudy (anonymous):

anita read 150 pages in 5 hours. what is her reading rate in pages per minute?

TheSmartOne (thesmartone):

If she read 150 pages in 5 hours then you just divide by 5 to find how many pages she reads in 1 hour. And 1 hour is 60 minutes so you divide agian by 60 to get it into minutes

TheSmartOne (thesmartone):

\[\frac{ 150~pages }{ 5~hours }=\frac{ 30~pgs }{ 1~hr }=\frac{ 30~pgs }{ 60~mins }=\frac{ .5~pgs }{ 1~mins }\]

TheSmartOne (thesmartone):

@yoyo5137